OpenStudy (anonymous):

Give a MEDAL who solve find how many nonnegative integer solutions of the system are here x+y<8 x+z<8 t+y<8 z+t<8

OpenStudy (caozeyuan):

try graphing all these regions on a single page on your calculator. FInd the intersection of these regions.

OpenStudy (caozeyuan):

Do you have a TI-NSpire CX

OpenStudy (caozeyuan):

I'll try, wait a minute

OpenStudy (caozeyuan):

sorry I was wrong, you have to graph it by hand, TI sucks on this one

OpenStudy (anonymous):

@amistre64 @Ashleyisakitty @ahoward79 @Falco276 @goformit100 @genius12

OpenStudy (amistre64):

4 equations with 4 unknowns

OpenStudy (amistre64):

x+y<8 ; y < 8-x x+z<8 ; z < 8-x well, y = z t+8-x<8 x = t

OpenStudy (amistre64):

8-x >= 0 , when x <= 8

OpenStudy (amistre64):

seems to me that the solution set is: x,y,z,t = (0,8]

OpenStudy (anonymous):

@amistre64 not obligatory y=z

OpenStudy (amistre64):

x+y < 8 0+7 ,1+6, 2+5, 3+4, 4+3, 5+2, 6+1 x + z < 8 0+7 ,1+6, 2+5, 3+4, 4+3, 5+2, 6+1 t+y<8 0+7 ,1+6, 2+5, 3+4, 4+3, 5+2, 6+1 z+t<8 0+7 ,1+6, 2+5, 3+4, 4+3, 5+2, 6+1

OpenStudy (amistre64):

forgot to tack onthe (7+0)s

OpenStudy (amistre64):

when x=0, y=7, z=7, t=0 when x=1, y=6, z=6, t=1

OpenStudy (amistre64):

at any rate, the set is the cross of {0,1,2,3,4,5,6,7}^4

OpenStudy (anonymous):

no when x+y<8 you also can have 1,2 1,3 1,4 1,5 1,0 1,1 1,6

OpenStudy (anonymous):

you are making x+y=7

OpenStudy (anonymous):

but the other cases?

OpenStudy (amistre64):

im simply defining a limit to it

OpenStudy (amistre64):

i beleive there are 7^4 elements of the set (x,y,z,t)

OpenStudy (amistre64):

or is that 8^4

OpenStudy (amistre64):

at any rate, its a lot of ordered pairs

OpenStudy (queelius):

This is a tricky counting problem. You could try to devise some case-by-case counting scheme, or you could enumerate all the possibilities (that satisfy the constraints) using a computer programming language. It's actually pretty straightforward if you know how to program. numSolutions = 0 for x in [0, 7] ..for y in [0, 7] ....for z in [0, 7] ......for t in [0, 7] ........if x + y < 8 and x + z < 8 and t + y < 8 and z + t < 8 then ..........numSolutions = numSolutions + 1 print numSolutions This will be extremely quick, since there are only 8^4 iterations.

OpenStudy (queelius):

Btw, it's 876.

OpenStudy (anonymous):

yes its corect